# Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1

### v3

@@ Appendix E Testing Process
The testing process for MolEdit task is demonstrated in Algorithm 1. However, in this process, we can not guarantee that the modified molecule differs the original molecule only in the specified functional group. The reasons are to follow:
  • The complexity of molecular structures makes it difficult to use regular matching to monitor the exact number of functional groups present. This is because molecules can have a wide variety of bonding patterns and spatial arrangements that are not always amenable to simple, rule-based matching techniques.
  • Modifying functional groups in specific positions can indeed lead to unintended changes in the molecule’s overall structure. This is due to the interconnected nature of atoms within a molecule, where altering one part can have a ripple effect on the stability and geometry of the entire structure.
  • The tools and algorithms currently available may not have the necessary sophistication to predict and control all the possible outcomes of a functional group modification, particularly when dealing with complex molecules or reactions that are not fully understood.
Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
@@
0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for atomatom in AA do         if |atom||atom|($m^{g}$) $\neq$ RR[atomatomthen            flag = false else if $t$ is BondNum then    for bondbond in $B$ do         if |bond||bond|($m^{g}$) $\neq$ RR[bondbondthen            flag = false else if $t$ is FunctionalGroup then    for groupgroup in $G$ do         if |group||group|($m^{g}$) $\neq$ RR[groupgroupthen            flag = false return flag
For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
Appendix F Detailed Results
In this section, we first show the leaderboard of S2-Bench in Table 12, where Claude-3.5 achieves the second place with an average weighted success rate of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama3.1-8B achieves the first place, which outperforms all the other LLMs.
Rank Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
1 Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33
2 Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
3 Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
4 GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
5 GPT-4o (OpenAI, 2023) - 49.08 32.29
6 Claude-3 (Anthropic, 2024a) - 46.14 30.47
7 Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22
8 Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
9 Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
10 Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
11 Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
12 GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
13 Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
14 Gemma3-12B (Team et al., 2025) 12 26.28 15.00
15 Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 7 25.07 14.61
16 Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
17 Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
18 chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
19 Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
20 Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
21 yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
22 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
23 BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
24 MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
25 Llama3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
26 MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
27 MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
28 Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
Table 12: Leaderboard of S2-Bench.
Figure 5: Performance Comparison across several representative LLMs on S2-Bench.
We also compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5. Among the proprietary models, it can be concluded that Claude-3.5 achieves a powerful performance in these subtasks, except for the DelComponent task and the BondNum task. In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5. More importantly, we can also observe that Llama3.1-8B (OpenMolIns-xlarge) outperforms most of the other LLMs across these nine subtasks, achieving the SOTA performance.
Meanwhile, these radar maps also imply the difficulties of the subtasks. For example, for majority of the LLMs, they are struggling with the BondNum task. Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
Finally, we also include a reasoning LLM, Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), as shown in Table 12, which achieves the best performance among 7 and 8B level open-source general LLMs, showing the potential of LLM reasoning in open-domain natural language-driven molecule generation tasks.
@@ Appendix F Detailed Results
Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15. Note that we do not provide results of vanilla galactica-125M, for it’s not instruction tuned and cannot respond to human instructions. All results of vanilla galactica-125M should be close to zeros.
@@
Claude-3.5 (Anthropic, 2024b) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
Claude-3 (Anthropic, 2024a) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
@@
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4336 0.5257 0.2279 0.5910 0.3401 0.6424 0.2185 0.5076
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
@@
Table 13: Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Claude-3.5 (Anthropic, 2024b) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
Claude-3 (Anthropic, 2024a) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
@@
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4336 0.5257 0.2279 0.5910 0.2655 0.4499 0.1194 0.6158
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
@@
Table 14: Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Claude-3.5 (Anthropic, 2024b) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
Claude-3 (Anthropic, 2024a) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
@@
Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
Appendix G Limitations
@@
Data Distribution. In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions. Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately.
The Selection of Subtasks. For the selection of subtasks, there are various choices in MolOpt, we compeletly understand that there are multiple viable options, especially when there are enormous metrics for evaluating molecular properties. However, we need to 1) consider the consistency between the three major tasks in S2-Bench and balance their weights (which means we can not select too many metrics for MolOpt, otherwise it would make the benchmark solely an optimization benchmark), and 2) ensure that the metrics can be efficiently and robustly calculated. Therefore, we only select the three common metrics, LogP, MR, and QRD, for MolOpt.
Appendix H Case Study
Finally, we include the case study to better illustrate the performance gap among different models.
Figure 6: Case study of AddComponent. In this example, we require LLMs to add a hydroxyl to the given molecule. Here, GPT-4o and Claude-3.5 barely modified the molecule and failed to add a hydroxyl to the given molecule, while other four models successfully passed the testing process. However, it can be clearly noted that although BioT5-base passed the testing process, the generated molecule of BioT5-base could be quite different from the original molecule.
Figure 7: Case study of DelComponent. In this example, we require LLMs to remove an amine from the given molecule. Here, GPT-4o, Claude-3.5, and BioT5-base all removed two amines, while the remaining models successfully passed the testing process.
Figure 8: Case study of SubComponent. In this example, we require LLMs to substitute the hydroxyl from the given molecule by a aldehyde. Here, although BioT5-base added a aldehyde, it failed to remove the hydroxyl. In contrast, the remaining models successfully passed the testing process.
Figure 9: Case study of LogP. In this example, we require LLMs to modfify the molecule to have a higher LogP value. Here, after the calculation of LogP values, we found all the generated molecules satisfy the requirement.
Figure 10: Case study of MR. In this example, we require LLMs to modify the molecule to have a higher MR value. Here, Llama3.1-8B-Instruct and Llama3.1-8B-Instruct (OpenMolIns-large) surprisingly failed the testing process. From the two errors, we could observe that the two models were trying to introduce extra rings to increase the MR value, while the ring numbers were misassigned.
Figure 11: Case study of QED. In this example, we require LLMs to lower the QED value of the given molecule. Compared to LogP and MR, QED is more difficult to predict. We could observe that GPT-4o, BioT5-base, Llama3.1-8B-Instruct, and Llama3.1-8B-Instruct (OpenMolIns-large) all increased the QED value of the molecule, while Claude-3.5 and Galactica-125M (OpenMolIns-xlarge) successfully captured the requirement and lowered the QED value.
Figure 12: Case study of AtomNum. In this example, we require LLMs to generate a molecule with 15 carbon atoms and 2 bromine atoms. Here, GPT-4o, Claude-3.5, BioT5-base, and Llama3.1-8B-Instruct all generated incorrect numbers of the specified atoms.
Figure 13: Case study of BondNum. In this example, we require LLMs to generate a molecule with 6 aromatic bonds, which normally exist in benzene rings. Here, GPT-4o, Claude-3.5, Llama3.1-8B-Instruct (OpenMolIns-large), and Galactica-125M (OpenMolIns-xlarge) captured the concept and generated molecules with a benzene ring in their structures. In contrast, BioT5-base failed to identify the requirement and Llama3.1-8B-Instruct generated too may aromatic rings.
Figure 14: Case study of FunctionalGroup. In this example, we require LLMs to generate a molecule with 2 benzene rings. Here, only Claude-3.5 failed to generate the correct number of benzene rings.

### v4

@@ Appendix A Testing Process
0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then for atomatom in AA do if |atom||atom|($m^{g}$) $\neq$ RR[atomatom] then flag = false else if $t$ is BondNum then for bondbond in $B$ do if |bond||bond|($m^{g}$) $\neq$ RR[bondbond] then flag = false else if $t$ is FunctionalGroup then for groupgroup in $G$ do if |group||group|($m^{g}$) $\neq$ RR[groupgroup] then flag = false return flag
Appendix B Detailed Results
@@
In this section, we show the leaderboard of S2-Bench in Table 8. The detailed experimental results of all the subtasks in Table 9, 10, and 11. Note that we do not provide results of vanilla galactica-125M, for it’s not instruction-tuned and cannot respond to human instructions. All results of vanilla galactica-125M should be close to zero.
Table 8. Leaderboard of S2-Bench.
Model # Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$ Rank
Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33 1
Claude-3.5 (Anthropic, 2024a) - 51.10 35.92 2
Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80 3
GPT-4-turbo (OpenAI, 2023) - 50.74 34.23 4
GPT-4o (OpenAI, 2023) - 49.08 32.29 5
Claude-3 (Anthropic, 2024b) - 46.14 30.47 6
Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22 7
Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73 8
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93 9
Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42 10
Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89 11
GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58 12
Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18 13
Gemma3-12B (Team et al., 2025) 12 26.28 15.00 14
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 7 25.07 14.61 15
Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.12 13.78 16
Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75 17
chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7) 18
Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6) 19
ChemDFM-v1.5-8B (Zhao et al., 2025) 8 18.24 12.07 20
ChemLLM-20B (Zhang et al., 2024) 20 16.23 9.76 21
SmileyLlama (Cavanagh et al., 2026) 8 16.18 8.43 22
Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10 23
yi-1.5-9B (Young et al., 2024) 9 14.10 7.32 24
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81 25
BioT5-base (Pei et al., 2023) 0.25 24.19 4.21 26
MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89 27
Llama3.2-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99 28
MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0) 29
MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9) 30
Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15 31
Table 9. Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Claude-3.5 (Anthropic, 2024a) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
Claude-3 (Anthropic, 2024b) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
@@
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4416 0.5199 0.2296 0.5832 0.3401 0.6424 0.2185 0.5076
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
SmileyLlama (Cavanagh et al., 2026) 0.1990 0.5460 0.1088 0.7700 0.0740 0.5990 0.0445 0.7940 0.0170 0.5870 0.0102 0.7740
ChemDFM-v1.5-8B (Zhao et al., 2025) 0.2942 0.6454 0.1899 0.9112 0.2056 0.6937 0.1426 0.9286 0.2198 0.6432 0.1414 0.8652
ChemLLM-20B (Zhang et al., 2024) 0.1436 0.5978 0.0858 0.2108 0.2092 0.5734 0.1200 0.3098 0.3088 0.6805 0.2101 0.4936
@@
Table 10. Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Claude-3.5 (Anthropic, 2024a) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
Claude-3 (Anthropic, 2024b) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
@@
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4134 0.4794 0.1982 0.5162 0.2655 0.4499 0.1194 0.6158
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
SmileyLlama (Cavanagh et al., 2026) 0.4520 0.4960 0.2241 0.7860 0.4070 0.5180 0.2111 0.7730 0.2900 0.5170 0.1499 0.7800
ChemDFM-v1.5-8B (Zhao et al., 2025) 0.2998 0.6613 0.1983 0.7582 0.2510 0.6831 0.1715 0.8170 0.3112 0.6600 0.2054 0.8072
ChemLLM-20B (Zhang et al., 2024) 0.2992 0.5634 0.1686 0.3788 0.1316 0.5452 0.0717 0.293 0.2594 0.5739 0.1489 0.3856
@@
Table 11. Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Claude-3.5 (Anthropic, 2024a) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
Claude-3 (Anthropic, 2024b) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
@@
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
SmileyLlama (Cavanagh et al., 2026) 0.0002 0.6560 0.0001 0.6420 0.0116 0.5970 0.0069 0.6990 0.0052 0.5970 0.0031 0.7430
ChemDFM-v1.5-8B (Zhao et al., 2025) 0.0112 0.6809 0.0076 0.6700 0.0172 0.6146 0.0106 0.8968 0.0312 0.6011 0.0188 0.8094
ChemLLM-20B (Zhang et al., 2024) 0.0288 0.7055 0.0203 0.3530 0.0204 0.6786 0.0138 0.3678 0.0600 0.6452 0.0387 0.5442
@@
Appendix C Limitations
@@
Data Distribution. Our allocation of atoms, bonds, and functional groups is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios.
Application to Complex Real-world Constraints. Incorporating constraints like stereochemistry, 3D geometry, ring/bridge topology, and basic synthesizability would represent more important aspects of real-world molecular design.
Realism of Instructions. Template-generated prompts may not fully capture the richness of real-world chemist queries.
Appendix D Broader Impacts
S2-Bench firstly proposes the text-based open molecule generation task for LLMs, aiming at adopting LLMs as chemist assistants and molecule operators for molecule discovery. Hopefully, our benchmark could inspire more research in this promising area, thereby accelerating the pace of drug discovery and material design. Ultimately, the development in molecule discovery will benefit human welfare in various ways. For example, the effective medicines could help save people’s lives and expand their lifespan.
Appendix E GenAI Disclosure
During the preparation of this work, the author(s) used LLMs to improve the language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.